The sum of two numbers is 30 and their difference is 10. What are the two numbers?
Answer by checkley77(12844)

X+Y=30
X-Y=10 ADD
---------------
2X=40
X=40/2
X=20 ANS.
20+Y=30
Y=30-20
Y=10 ANS.
PROOF:
20-10=10
10=10